# Multi-Agent Autonomous Driving Systems with Large Language Models: A Survey of Recent Advances

### v1

@@ 4.1 Instructor Paradigm
In Figure 5, the instructor paradigm involves agents interacting with humans in a conversational manner, where humans act as “tutors" to offer quantitative and qualitative feedback to improve the agent’s decision-making Li et al. (2017). Quantitative feedback typically includes binary evaluations or ratings, while qualitative feedback consists of language suggestions for refinement. The agent incorporates these suggestions to adapt and enhance its performance in complex driving scenarios. For instance, Wang et al. (2023) propose “Expert-Oriented Black-box Tuning", a method where domain experts provide feedback to optimize model performance. Similarly, Ma et al. (2024) present a human-guided learning pipeline that integrates driver feedback to refine agent decision-making.
@@ 4.2 Partnership Paradigm
In Figure 5, the partnership paradigm emphasizes collaboration, where agents and humans interact as equals to accomplish complex driving tasks. In this paradigm, agents assist in decision-making by adapting to individual driver preferences and real-time traffic conditions. For instance, Talk2Drive Cui et al. (2023), DaYS Cui et al. (2024a) and Receive Cui et al. (2024b) utilize memory modules to store human-vehicle interactions, enabling a more personalized driving experience based on individual driver preferences, such as overtaking speed and following distance. Additionally, infrastructure agents in AccidentGPT Wang et al. (2024b) and ConnectGPT Tong and Solmaz (2024) connect vehicles to monitor traffic conditions, identify potential hazards, and provide proactive safety warnings, blind spot alerts, and driving suggestions through agent-human interaction.
@@ 5.2 Collaborative Decision-Making
LanguageMPC Sha et al. (2023) uses LLMs to perform scenario analysis and decision-making. Additionally, it introduces a multi-vehicle control method where distributed LLMs govern individual vehicle operations, while a central LLM facilitates multi-vehicle communication and coordination. AgentsCoDriver Hu et al. (2024a) presents a comprehensive LLM-based multi-vehicle collaborative decision-making framework with life-long learning capabilities, moving the field towards practical applications. This framework consists of five parts, as follows: the observation module, cognitive memory module, and reasoning engine support the high-level decision-making process for AD; the communication module enables negotiation and collaboration among vehicles; and the reinforcement reflection module reflects the output and decision-making process. Similarly, AgentsCoMerge Hu et al. (2024b) combines vision-based and text-based scene understanding to gather essential environmental information and incorporates a hierarchical planning module to allow agents to make informed decisions and effectively plan trajectories. Instead of directly interacting with each other, agents in KoMA Jiang et al. (2024) analyze and infer the intentions of surrounding vehicles via an interaction module to enhance decision-making. It also introduces a shared memory module to store successful driving experiences and a ranking-based reflection module to review them.
5.3 Other Collaborative Assistance-Tools
The long-term data accumulation in both industry and academia has enabled great success in highway driving and automatic parking Liu et al. (2024b). However, collecting real-world data remains costly, especially for multi-agents or customized scenarios. Additionally, the uncontrollable nature of real scenarios makes it challenging to capture certain corner cases. To address these issues, many LLM-based studies focus on simulating multi-agent ADS, offering a cost-effective alternative to real-world data collection. For example, ChatSim Wei et al. (2024) provides editable photo-realistic 3D driving scenario simulations via natural language commands and external digital assets. The system leverages multiple LLM agents with specialized roles to decompose complex commands into specific editing tasks, introducing novel McNeRF and Mclight methods that generate customized high-quality output. HumanSim Zhou et al. (2024a) integrates LLMs to simulate human-like driving behaviors in multi-agent systems via pre-defined driver characters. By employing navigation strategies, HumanSim facilitates behavior-level control of vehicle movements, making it easier to generate corner cases in multi-agent environments.
@@ 5.3 Other Collaborative Assistance-Tools
Although many innovative studies have explored the application of LLM-based multi-agent ADS, significant technical challenges remain in deploying LLMs locally on autonomous vehicles due to their huge computational resource requirements Sun et al. (2024). To address these issues, Tang et al. (2024) apply remote LLMs to provide assistance for connected autonomous vehicles, which communicate between themselves and with LLMs via vehicle-to-everything technologies. Moreover, this study evaluates LLMs’ comprehension of driving theory and skills in a manner akin to human driver tests. However, remote LLM deployment can introduce inference latency, posing risks in emergency scenarios. To further improve system efficiency, Chen et al. (2024a) introduce a novel edge-cloud collaborative ADS with drift detection capabilities, using small LLMs on edge devices and GPT-4 on cloud to process motion planning data and complex inference tasks, respectively.
In addition, ALGPT Zhou et al. (2024c) uses a muti-agent cooperative framework to enable open-vocabulary and multimodal auto-annotation for autonomous driving. ALGPT introduces a Standard Operating Procedure that clarifies the role of each agent and shares project documentation, thereby enhancing the effectiveness of multi-agent interactions. Moreover, ALGPT establishes a specialized knowledge base for each type of agent, using CoT and In-Context Learning Brown et al. (2020).
6 Datasets
We organize the latest state-of-the-art open-source work to foster research of more advanced ADSs. And we summarize mainstream ADS datasets in Table 1. More details are listed in Appendix A.3.
@@ 7 Challenges and Future Directions
  • Hallucination Problem. It refers to LLMs generating outputs that are factually incorrect or non-sensical Huang et al. (2023). In complex driving scenarios, a single driving agent’s hallucinations in an LLM-based multi-agent ADS can be accepted and further propagated by other agents in the network via the inter-agent communication, potentially leading to serious accidents. Consequently, detecting and mitigating hallucinations at the individual agent level and managing the flow of information between agents are crucial issues for future research Fan et al. (2024).
  • Multi-Modality Ability. Agents in current multi-agent systems primarily use LLMs for scene understanding and decision making. These methods convert the outputs of perception algorithms into textual representations through manual prompts or interpreters, which are then fed into an LLM to produce decisions. This approach heavily depends on the performance of the perception algorithm and can lead to loss of environmental information Gao et al. (2023). Therefore, integrating language understanding with the ability to process and fuse multiple data modalities to develop a multimodal multi-agent ADS represents a promising direction for future research.
  • Scalability Problem. LLM-based multi-agent ADS can scale up by adding more agents to handle increasingly complex driving scenarios. However, more LLM agents increase the demand for computing resources, while their interactions impose strict requirements on communication efficiency, which is critical for real-time decision-making Huang et al. (2024b). Therefore, under limited computing resources, it is crucial to develop a system architecture that supports distributed computing and efficient communication, as well as agents capable of adapting to various environments and tasks, to optimize multi-agent ADS within resource constraints.
@@ 8 Conclusion
This paper systematically outlines LLM-based multi-agent ADS and comprehensively reviews the latest research in this field. Our study first traces the development trajectory of LLM-based multi-agent ADS from single-agent ADS to multi-agent ADS. Subsequently, we provide a detailed description of the LLM-based multi-agent ADS from the perspectives of agent-environments and profiles, inter-agent interaction mechanisms, and agent-human interactions. We also systematically classify and introduce existing studies from the perspectives of multi-agent interaction, agent-human interaction, and different applications. Finally, this paper provides comprehensive public datasets and open source codes, and deeply explores the current challenges and future research directions of LLM-based multi-agent ADS. We hope that this review can bring new inspiration and ideas to future research on LLM-based multi-agent ADS.
@@
Emerging Research and Limited Data. As the field of LLM-based multi-agent ADS is still relatively new, existing research is limited, which may restrict the scope of our classification and analysis. Some Unverified Work. Since LLM-based multi-agent ADS is a novel topic, some of the papers summarized in this review are from unreviewed arXiv preprints. As these works have not been formally published, their conclusions may require further investigation to confirm their validity. Limited Discussion on Real-world Applications. As LLM-based multi-agent ADS is still in the theoretical stage, although many companies have begun deploying practical applications in this area, this review does not cover discussions on real-world deployments due to a lack of up-to-date internal information from these companies.
Appendix A Appendix
A.1 Data-driven Autonomous Driving System
@@
Traditional ADS rely on data-driven approaches, which are categorized into modular and end-to-end frameworks Chen et al. (2024b). Modular-based systems break the entire autonomous driving process into separate components, such as perception module, prediction module, and planning module. Perception modules are responsible for obtaining information about the vehicle’s surrounding environment, aiming to identify and locate important traffic elements such as obstacles, pedestrians, and vehicles near the autonomous vehicle, usually including tasks such as object detection Wang et al. (2021) and object occupancy prediction Tong et al. (2023). Prediction modules estimate the future motions of surrounding traffic participants based on the information provided by the perception module, usually including tasks such as trajectory prediction and motion prediction Shi et al. (2022). Planning module aims to derive safe and comfortable driving routes and decisions through the results of perception and prediction Sauer et al. (2018). Each module is individually developed and integrated into onboard vehicles to achieve safe and efficient autonomous driving functions. Although modular methods have achieved remarkable results in many driving scenarios, the stacking design of multiple modules can lead to the loss of key information during transmission and introduce redundant calculations. Furthermore, due to the inconsistency in the optimization objectives of each module, the modular-based system may accumulate errors, which can negatively impact the vehicle’s overall decision-making performance. End-to-end-based systems integrate the entire driving process into a single neural network, and then directly optimize the entire driving pipeline from sensor inputs to produce driving actions Chen et al. (2024b). However, this approach introduces the “black box” problem, meaning a lack of transparency in the decision-making process, complicating interpretation and validation.
A.2 LLMs in Autonomous Driving System
@@
As shown in Figure 6, 7, LLMs, with their powerful open-world cognitive and reasoning capabilities, have shown significant potential in ADSs Yang et al. (2023); Li et al. (2023). LC-LLM Peng et al. (2024) is an explainable lane change prediction model that leverages LLMs to process driving scenario information as natural language prompts. By incorporating CoT reasoning and supervised finetuning, it not only predicts lane change intentions and trajectories but also provides transparent and reliable explanations for its predictions. GPT-Driver Mao et al. (2023) regards the motion planning task as a language modeling problem, using a fine-tuned GPT-3.5 model Ye et al. (2023) to generate driving trajectories. DriveGPT4 Xu et al. (2024) introduces an interpretable end-to-end autonomous driving system that uses multimodal LLMs to process multi-frame video inputs and textual queries, enabling vehicle action interpretation and low-level control prediction. By employing a visual instruction tuning dataset and mixfinetuning strategy, it provides a novel approach to directly map sensory inputs to actions, achieving superior performance in autonomous driving tasks. Driving with LLM Chen et al. (2024c) integrates vectorized numeric data with pre-trained LLMs to improve context understanding in driving scenarios and enhances the interpretability of driving decisions.
@@
Figure 7: The communication among multiple agents in an LLM-based multi-agent system Hu et al. (2024a)
A.3 Datasets
Single-agent Autonomous Driving Dataset. Single-agent datasets are obtained from a single reference agent, which can be the ego vehicle or roadside infrastructure, using various sensors. Mainstream singel-agent autonomous driving datasets like KITTI Geiger et al. (2012), nuScenes Geiger et al. (2020), and Waymo Sun et al. (2020) provide comprehensive multimodal sensor data, enabling researchers to develop and benchmark algorithms for multiple tasks such as object detection, tracking, and segmentation.
In addition to these foundational datasets, newer ones like BDD-X Kim et al. (2018), DriveLM Sima et al. (2025), and nuScenes-QA Qian et al. (2024) introduce action descriptions, detailed captions, and question-answer pairs that can be used to interact with LLMs. Combining language information with visual data can enrich semantic and contextual understanding, promote a deeper understanding of driving scenarios, and enhance the safety and interaction capabilities of autonomous vehicles.

### v2

@@ 4.1 Instructor Paradigm
In Figure 5, the instructor paradigm involves agents interacting with humans in a conversational manner, where humans act as “tutors” to offer quantitative and qualitative feedback to improve agent decision-making Li et al. (2017). Quantitative feedback typically includes binary evaluations or ratings, while qualitative feedback consists of language suggestions for refinement. Agents incorporate this feedback to adapt and perform better in complex driving scenarios. For instance, Wang et al. (2023) propose “Expert-Oriented Black-box Tuning”, where domain experts provide feedback to optimize model performance. Similarly, Ma et al. (2024) present a human-guided learning pipeline that integrates driver feedback to refine agent decision-making.
@@ 4.2 Partnership Paradigm
As shown in Figure 5, the partnership paradigm emphasizes collaboration, where agents and humans interact as equals to accomplish complex driving tasks. In this paradigm, agents assist in decision-making by adapting to individual driver preferences and real-time traffic conditions. For instance, Talk2Drive Cui et al. (2023), DaYS Cui et al. (2024a) and Receive Cui et al. (2024b) utilize memory modules to store human-vehicle interactions, enabling a more personalized driving experience based on individual driver preferences, such as overtaking speed and following distance. Additionally, infrastructure agents in AccidentGPT Wang et al. (2024b) and ConnectGPT Tong and Solmaz (2024) connect vehicles to monitor traffic conditions, identify potential hazards, and provide proactive safety warnings, blind spot alerts, and driving suggestions through agent-human interaction.
@@ 5.2 Collaborative Decision-Making
LanguageMPC Sha et al. (2023) uses LLMs to perform scenario analysis and decision-making. Additionally, it introduces a multi-vehicle control method where distributed LLMs govern individual vehicle operations, while a central LLM facilitates multi-vehicle communication and coordination. AgentsCoDriver Hu et al. (2024a) presents a comprehensive LLM-based multi-vehicle collaborative decision-making framework with life-long learning capabilities, moving the field towards practical applications. This framework consists of five parts, as follows: the observation module, cognitive memory module, and reasoning engine support the high-level decision-making process for AD; the communication module enables negotiation and collaboration among vehicles; and the reinforcement reflection module reflects the output and decision-making process. Similarly, AgentsCoMerge Hu et al. (2024b) combines vision-based and text-based scene understanding to gather essential environmental information and incorporates a hierarchical planning module to allow agents to make informed decisions and effectively plan trajectories. Instead of directly interacting with each other, agents in KoMA Jiang et al. (2024a) analyze and infer the intentions of surrounding vehicles via an interaction module to enhance decision-making. It also introduces a shared memory module to store successful driving experiences and a ranking-based reflection module to review them.
5.3 Collaborative Cloud-Edge Deployment
Although many innovative studies have explored the application of LLM-based multi-agent ADS, significant technical challenges remain in deploying LLMs locally on autonomous vehicles due to their huge computational resource requirements Sun et al. (2024a). To address these issues, Tang et al. (2024) apply remote LLMs to provide assistance for connected autonomous vehicles, which communicate between themselves and with LLMs via vehicle-to-everything technologies. Moreover, this study evaluates LLMs’ comprehension of driving theory and skills in a manner akin to human driver tests. However, remote LLM deployment can introduce inference latency, posing risks in emergency scenarios. To further improve system efficiency, Chen et al. (2024a) introduce a novel edge-cloud collaborative ADS with drift detection capabilities, using small LLMs on edge devices and GPT-4 on cloud to process motion planning data and complex inference tasks, respectively.
@@ 5.3 Collaborative Cloud-Edge Deployment
5.4 Collaborative Assistance-Tools
The long-term data accumulation in both industry and academia has enabled great success in highway driving and automatic parking Liu et al. (2024b). However, collecting real-world data remains costly, especially for multi-agents or customized scenarios. Additionally, the uncontrollable nature of real scenarios makes it challenging to capture certain corner cases. To address these issues, many LLM-based studies focus on simulating multi-agent ADS, offering a cost-effective alternative to real-world data collection. For example, ChatSim Wei et al. (2024) provides editable photo-realistic 3D driving scenario simulations via natural language commands and external digital assets. The system leverages multiple LLM agents with specialized roles to decompose complex commands into specific editing tasks, introducing novel McNeRF and Mclight methods that generate customized high-quality output. HumanSim Zhou et al. (2024a) integrates LLMs to simulate human-like driving behaviors in multi-agent systems via pre-defined driver characters. By employing navigation strategies, HumanSim facilitates behavior-level control of vehicle movements, making it easier to generate corner cases in multi-agent environments. In addition, ALGPT Zhou et al. (2024c) uses a multiagent cooperative framework for open-vocabulary, multimodal auto-annotation in autonomous driving. It introduces a Standard Operating Procedure to define agent roles and share documentation, enhancing interaction effectiveness. ALGPT also builds specialized knowledge bases for each agent using CoT and In-Context Learning Brown et al. (2020).
6 Datasets and Benchmark
We organize recent open-source work to foster research on advanced ADSs. Mainstream ADS datasets are summarized in Table 2.
Single-Agent Autonomous Driving Data. Single-agent datasets are obtained from a single reference agent, which can be the ego vehicle or roadside infrastructure, using various sensors. Mainstream singel-agent autonomous driving datasets like KITTI Geiger et al. (2012), nuScenes Geiger et al. (2020), and Waymo Sun et al. (2020) provide comprehensive multimodal sensor data, enabling researchers to develop and benchmark algorithms for multiple tasks such as object detection, object tracking, and object segmentation. In addition to these foundational datasets, newer ones like BDD-X Kim et al. (2018), DriveLM Sima et al. (2025), and nuScenes-QA Qian et al. (2024) introduce action descriptions, detailed captions, and question-answer pairs that can be used to interact with LLMs. Combining language information with visual data can enrich semantic and contextual understanding, promote a deeper understanding of driving scenarios, and improve the safety and interaction capabilities of autonomous vehicles.
Benchmarks. Several benchmarks are particularly well-suited for evaluating collaborative decision-making in autonomous driving. The INTERACTION dataset (Zhan et al., 2019) includes a variety of real-world interactive scenarios, such as roundabouts and lane merging. It provides vehicle trajectories that enable an assessment of cooperative maneuvering and negotiation behaviors. Another important benchmark is the Waymo Open Motion Dataset (Ettinger et al., 2021), which is explicitly designed for interactive multi-agent motion prediction and planning. It features challenging scenarios, including merges and unprotected left turns, along with detailed annotations of interactive agents. In addition, the SMARTS benchmark (Zhou et al., 2021) offers standardized scenarios for multi-agent autonomous driving research, particularly focusing on ramp merging and navigating unsignalized intersections. This work allows for direct comparisons of algorithms in cooperative traffic management tasks. These benchmarks provide comprehensive test bases for evaluating the coordination, safety, and adaptability of LLM-based multi-agent ADSs.
@@ 7 Challenges and Future Directions
❶ Hallucination, Safety & Trustworthiness. Hallucination refers to LLMs generating outputs that are factually incorrect or non-sensical Huang et al. (2023). In complex driving scenarios, a single driving agent’s hallucinations in an LLM-based multi-agent ADS can be accepted and further propagated by other agents in the network via the inter-agent communication, potentially leading to serious accidents. Detecting agent-level hallucinations and managing inter-agent information flow are key to improving system safety and trust Fan et al. (2024). Recent advances in spatiotemporal traffic analysis Zhang et al. (2024d); Jiang et al. (2024b) further support real-time condition assessment, improving vehicle-road interaction and overall safety of ADS.
❷ Legal, Security & Privacy. As agents autonomously exchange and process information within multi-agent ADS, the distribution of legal liability between individual users and manufacturers becomes ambiguous, particularly in cases involving system failures or collisions. In addition, vulnerable communication methods and strict user privacy requirements place high demands on cryptographic protocols and data management. These interrelated concerns collectively represent critical directions for future research and regulatory initiatives.
❸ Multi-Modality Ability. In current multi-agent systems, agents primarily use LLMs for scene understanding and decision-making. Perception outputs are converted into text via manual prompts or interpreters, then processed by LLMs to generate decisions. This pipeline is limited by perception performance and may cause information loss Gao et al. (2023). Integrating language understanding with multimodal data fusion offers a promising direction for future multimodal multi-agent ADSs.
❹ Real-World Deployment & Scalability. LLM-based multi-agent ADS can scale up by adding more agents to handle increasingly complex driving scenarios. However, more LLM agents increase the demand for computing resources, while their interactions impose strict requirements on communication efficiency, which is critical for real-time decision-making Huang et al. (2024b). Therefore, under limited computing resources, it is crucial to develop a system architecture that supports distributed computing and efficient communication, as well as agents capable of adapting to various real-world environments and tasks, to optimize multi-agent ADS within resource constraints.
❺ Human-Agent Interaction. Current multi-agent ADS struggle to communicate intentions to human road users, relying on static signals inadequate for complex scenarios. Developing LLM-powered adaptive interfaces that generate context-appropriate, human-understandable communications while maintaining safety and trust presents a key deployment challenge Xia et al. (2025).
@@ 8 Conclusion
This paper systematically reviews LLM-based multi-agent ADSs and traces their evolution from single-agent to multi-agent systems. We detail their core components, including agent environments and profiles, inter-agent interaction, and agent-human communication. Existing studies are categorized by interaction types and applications. We further compile public datasets and open-source implementations, and discuss challenges and future directions. We hope this review will inspire NLP community to explore more practical and impactful applications in LLM-based multi-agent ADS.
@@
Despite being a survey, this work still has several limitations. ❶ Emerging Research and Limited Data. As LLM-based multi-agent ADS is an emerging field, the current body of research is still growing. While this may limit the breadth of our classification, we have aimed to provide a representative and forward-looking overview based on the most relevant and recent work. ❷ Some Unverified Work. Given the novelty of this topic, some referenced works are from unreviewed arXiv preprints. We include them to reflect the latest progress and ideas, while acknowledging that their findings may require further validation through peer review. ❸ Limited Discussion on Real-world Applications. Although industrial adoption of LLM-based multi-agent ADS is underway, public documentation remains limited. As a result, this review focuses on academic contributions, and real-world deployments are left for future investigation.
Appendix A Data-driven Autonomous Driving System
@@
Traditional ADS rely on data-driven approaches, which are categorized into modular and end-to-end frameworks Chen et al. (2024b). Modular-based systems break the entire autonomous driving process into separate components, such as perception module, prediction module, and planning module. Perception modules are responsible for obtaining information about the vehicle’s surrounding environment, aiming to identify and locate important traffic elements such as obstacles, pedestrians, and vehicles near the autonomous vehicle, usually including tasks such as object detection Wang et al. (2021) and object occupancy prediction Tong et al. (2023). Prediction modules estimate the future motions of surrounding traffic participants based on the information provided by the perception module, usually including tasks such as trajectory prediction and motion prediction Shi et al. (2022). Planning module aims to derive safe and comfortable driving routes and decisions through the results of perception and prediction Sauer et al. (2018). Each module is individually developed and integrated into onboard vehicles to achieve safe and efficient autonomous driving functions. Although modular methods have achieved remarkable results in many driving scenarios, the stacking design of multiple modules can lead to the loss of key information during transmission and introduce redundant calculations. Furthermore, due to the inconsistency in the optimization objectives of each module, the modular-based system may accumulate errors, which can negatively impact the vehicle’s overall decision-making performance. End-to-end-based systems integrate the entire driving process into a single neural network, and then directly optimize the entire driving pipeline from sensor inputs to produce driving actions Chen et al. (2024b). However, this method introduces the “black box” problem, meaning a lack of transparency in the decision-making process, complicating interpretation.
Appendix B LLMs in Autonomous Driving System
@@
As shown in Figure 6 and Figure 7, LLMs, with their powerful open-world cognitive and reasoning capabilities, have shown significant potential in ADSs Yang et al. (2023); Li et al. (2023). LC-LLM Peng et al. (2024) is an explainable lane change prediction model that leverages LLMs to process driving scenario information as natural language prompts. By incorporating CoT reasoning and supervised finetuning, it not only predicts lane change intentions and trajectories but also provides transparent and reliable explanations for its predictions. GPT-Driver Mao et al. (2023) regards the motion planning task as a language modeling problem, using a fine-tuned GPT-3.5 model Ye et al. (2023) to generate driving trajectories. DriveGPT4 Xu et al. (2024) introduces an interpretable end-to-end autonomous driving system that uses multimodal LLMs to process multi-frame video inputs and textual queries, enabling vehicle action interpretation and low-level control prediction. By employing a visual instruction tuning dataset and mixfinetuning strategy, it provides a novel approach to directly map sensory inputs to actions, achieving superior performance in autonomous driving tasks. Driving with LLM Chen et al. (2024c) integrates vectorized numeric data with pre-trained LLMs to improve context understanding in driving scenarios and enhances the interpretability of driving decisions.
@@
Figure 7: The communication among multiple agents in an LLM-based multi-agent system Hu et al. (2024a).
Appendix C LLM-enhanced Multi-Agent ADSs
To highlight the application of LLMs and other NLP technologies in multi-agent ADSs, we have specially prepared Table 4. This table systematically sorts out existing research from the two dimensions of “environment & subject characteristics” and “interaction mode”, and marks the LLMs used in each solution one by one. Our goal is to help readers quickly grasp the landscape of this cross-domain research and better understand how LLM capabilities are being adapted to complex ADS scenarios.
Appendix D Real-World Multi-Agent LLM Systems in Autonomous Driving
NVIDIA DriveOS LLM Integrationhttps://developer.nvidia.com/blog/streamline-llm-deployment-for-autonomous-vehicle-applications-with-nvidia-driveos-llm-sdk/ NVIDIA has released the DriveOS LLM SDK, which allows multiple AI agents (for perception, planning, user interaction) to run on edge computing, with the ability to infer LLMs. For example, a car can have a local LLM-based agent for real-time driving decisions or V2X message interpretation, offloading heavy computational tasks to optimized hardware. This lightweight onboard agent works in tandem with powerful cloud-based AI. In this conceptual multi-agent setup, the onboard LLM can handle immediate tasks and natural language commands while querying the cloud-based LLM for complex planning or traffic knowledge.
LLM in the Cabin and Beyondhttps://www.theverge.com/2023/6/16/23763208/mercedes-benz-chatgpt-voice-assistant-beta-test Mercedes-Benz is collaborating with NVIDIA to develop multimodal LLMs that can interpret sensor data and driver preferences to assist in driving decisions. LLM-based intelligent agents can act as “digital co-pilots”, monitoring the surrounding environment and providing maneuvering recommendations.
Table 3: Comparison of Interaction Modes and System Structures in LLM-Based Multi-Agent ADSs.
[…]
Table 4: Comparative Summary of LLM-Based Multi-Agent ADS Research.
[…]
LanguageMPC
(Sha et al., 2023)
2023/10 Simulation Pre-defined Vehicle agents, Human Centralized Cooperative Instructor Paradigm GPT-3.5
AgentsCoDriver
(Hu et al., 2024a)
2024/04 Simulation Pre-defined Vehicle agents Decentralized Cooperative - GPT-3.5-turbo
KoMA
(Jiang et al., 2024a)
2024/07 Simulation Pre-defined Vehicle agents Shared Message pool Cooperative Instructor Paradigm GPT-4
AgentsCoMerge
(Hu et al., 2024b)
2024/08 Simulation Pre-defined Vehicle agents Decentralized, Hierarchical Cooperative Instructor Paradigm GPT/Claude/Gemini Series
CoDrivingLLM
(Fang et al., 2024)
2024/09 Simulation Pre-defined Vehicle agents Centralized Cooperative Instructor Paradigm GPT-4o
CoMAL
(Yao et al., 2024)
2024/10 Simulation Pre-defined Vehicle agents, Human Shared Message pool Cooperative Instructor Paradigm GPT-4o-mini, Qwen-72B, Qwen-32B, Qwen-7B
Complement-Vehicle’s-FOV
(Dona et al., 2024)
2024/08 Simulation Pre-defined Vehicle agents, Infrastructure agents, Human Decentralized, Hierarchical, Centralized Cooperative Instructor Paradigm, Partnership Paradigm GPT-4V, GPT-4o
CAV-LLM-Driving-Assistant
(Tang et al., 2024)
2024/11 Simulation Pre-defined Vehicle agents,Human Decentralized Cooperative Instructor Paradigm GPT-4V, GPT-4o
EC-Drive
(Chen et al., 2024a)
2024/08 Simulation Pre-defined Vehicle agents, Infrastructure agents Hierarchical Cooperative Instructor Paradigm LLaMA-Adapter (7B), GPT-4
ChatSim
(Wei et al., 2024)
2024/02 Simulation Pre-defined, Model-generated Human, Assistant agents Hierarchical, Centralized Cooperative Instructor Paradigm GPT-4
ALGPT
(Zhou et al., 2024c)
2024/01 Simulation Pre-defined, Model-generated Assistant agents Hierarchical Cooperative - GPT series
AD-H
(Zhang et al., 2024c)
2024/06 Simulation Pre-defined Vehicle agents, Human Hierarchical Cooperative Instructor Paradigm LLaVA-7B-V1.5
SurrealDriver
(Jin et al., 2024)
2023/09 Simulation Pre-defined Vehicle agents, Infrastructure agents, Human Hierarchical Cooperative Instructor Paradigm GPT series, Llama, PaLM
LDPD
(Liu et al., 2024a)
2024/10 Simulation Model-generated Vehicle agents Hierarchical, Centralized Cooperative - GPT-4o
V-HOI MLCR
(Zhang et al., 2024a)
2024/03 Simulation Pre-defined Vehicle agents, Human Hierarchical Cooperative, Debate Instructor Paradigm GPT-4, GPT-3.5
Co-Pilot
(Wang et al., 2023)
2023 Physics Pre-defined Vehicle agents, Human Decentralized Cooperative Instructor Paradigm GPT-3.5-turbo-0301
PPE
(Ma et al., 2024)
2024 Simulation Pre-defined Vehicle agents, Human Decentralized Cooperative Partnership Paradigm GPT-4-turbo-preview and GPT-3.5-turbo
Drive-as-You-Speak
(Cui et al., 2024a)
2023/09 Simulation Pre-defined Vehicle agents, Human Decentralized Cooperative Instructor Paradigm, Partnership Paradigm GPT-4
Reason-and-React
(Cui et al., 2024b)
2023/10 Simulation Pre-defined Vehicle agents, Human Decentralized Cooperative Instructor Paradigm, Partnership Paradigm GPT-4
DriVLMe
(Huang et al., 2024a)
2024/06 Simulation Pre-defined Vehicle agents, Human Decentralized Cooperative Instructor Paradigm Vicuna-7B + LoRA
AccidentGPT
(Wang et al., 2024b)
2024/06 Physics Pre-defined Vehicle agents, Infrastructure agents, Human Hierarchical, Centralized, Decentralized Cooperative Instructor Paradigm GPT-4
ConnectGPT
(Tong and Solmaz, 2024)
2024/06 Physics Pre-defined Vehicle agents, Infrastructure agents, Human Hierarchical, Centralized, Decentralized Cooperative Instructor Paradigm GPT-4
DriveAgent
(Hou et al., 2025)
2025/05 Physics Pre-defined Assistant agents Decentralized Cooperative - LLaMA-3.2-Vision
CCMA
(Zhang et al., 2025)
2025 Simulation Pre-defined Vehicle agents, Assistant agents Hierarchical, Decentralized Cooperative - GLM-4v-9B
V2V-LLM
(Chiu et al., 2025)
2025/02 Simulation Pre-defined Vehicle agents Decentralized Cooperative - LLaVA-v1.5-7b
IITI
(Fang et al., 2025)
2025/03 Simulation Pre-defined Vehicle agents, Human Decentralized Cooperative Instructor Paradigm Llama3
Tell-drive
(Xu et al., 2025)
2025/02 Simulation Pre-defined Vehicle agents Hierarchical, Decentralized Cooperative - GPT-4o-min
Human-RLHF
(Sun et al., 2024b)
2024/06 Simulation Pre-defined Vehicle agents, Human Decentralized Cooperative Instructor Paradigm GPT-4o
GameChat
(Mahadevan et al., 2025)
2025/03 Simulation Pre-defined Vehicle agents, Human Decentralized Cooperative Instructor Paradigm GPT-4o-mini
hybrid LLM-DDQN
(Yan et al., 2025)
2024/10 Simulation Pre-defined Vehicle agents, Infrastructure agents Decentralized, Hierarchical Cooperative - GPT-3.5, Llama3.1-8B, Llama3.1-70B